A merchant has coffee worth $30 a pound that she wishes to mix with 90 pounds of coffee work $60 a pound to get a mixture that can be sold for $40 a pound. How many pounds of the $30 coffee should be used?
Answer by ptaylor(1894)

Let x=amount of the $30 coffee needed
($ understood)
Now we know that the value of the $30 coffee (30x) plus the value of the $60 coffee (60*90) has to equal the value of the final mixture(40* (90+x)). So our equation to solve is:
30x+60*90=40(90+x) get rid of parens and simplify
30x+5400=3600+40x subtract 30x and also 3600 from each side
30x-30x+5400-3600=3600-3600+40x-30x collect like terms
1800=10x divide each side by 10
x=180 lbs -------------amount of $30 coffee needed
CK
30*180+60*90=40*270
5400+5400=10800
10800=10800
Hope this helps---ptaylor